OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)?

OpenStudy (anonymous):

james I watched the video still struggleing with it

OpenStudy (anonymous):

6/-2= -3 y=-3x+3

OpenStudy (jamesj):

We all struggle with new concepts and need to review things a few times. So I recommend you watch it again.

OpenStudy (anonymous):

3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3 is this correct

OpenStudy (anonymous):

x+3y=3

OpenStudy (jamesj):

first, what is the slope of the line passing between those two points?

OpenStudy (anonymous):

6/2=3

OpenStudy (jamesj):

No. Draw a sketch with these two points and you'll see that the line between them must have negative slope, so that answer of +3 can't be right.